how do i get 10x^2+9=x into y=ax^2+bx+c to use the quadratic formula?
Answer by Alan3354(69443)

how do i get 10x^2+9=x into y=ax^2+bx+c to use the quadratic formula?
---------
10x^2+9=x
10x^2 - x + 9 = 0
Solve for x. y is not needed.